Reports 1-1 of 1 Clear search Modify search
Virgo Runs (O3)
berni - 15:02 Tuesday 11 June 2019 (46072) Print this report
operator report - morning shift
ITF found unlocked with WI and WE local controls open; Andrea was trying to close the WE loop but he had some difficulties.
I tried with the WI LC but I also had some difficulties (charged filter?) thus at 5:05 UTC I contacted the SUSP on call; he was able to close the loops from remote at 5:40 UTC.
In that period the ITF was set in Troubleshooting mode.

After that I tried to relock the ITF but it was very unstable and it was not possible to reach LOW_NOISE_1 collecting many unlocks at LOCKING_SSFS; at 6:15 UTC I stopped the locking because the Maintenance activities prevented the lock.

At 5:52 UTC started the planned Maintenance, here the activities communicated to the control loop:
- Fabozzi and external firm at NE;
- grass cutting;
- Vacuum activities with IB and DET Large valve closed;
- INJ activities during the maintenance ;
- operations on the NE Tiltmeter;
- cleaning operations;
- ENV activities at NE building, and .
- visitors in CB from 9:30 UTC to 10:15 UTC.
all the maintenance activities concluded at around 10:16 UTC.

After that we started the relock but the ITF was still unstable; Julia and Camilla started working on it.

At 11:19 UTC while trying to relock we had a problem with the SNEB_LC timing error; I set Troubleshooting and I called Alain who fixed the problem from remote. To be noted that at the same time Irene was plugging a magnetometer (see attached plot1).

At 12:17 UTC we could reach for the first time in the shift LOW_NOISE_1; at 12:41 UTC the ITF was locked in LOW_NOISE_3_SQZ but it was left in LOCKED because the ISYS team was working in EE room, see below.

ISYS
At 12:24 UTC the temperature of the Laser main chiller started to increase (red flag on DMS, plot2) and I promptly contacted the ISYS on-call. All the ISYS team went in EE room to investigate at the problem; still in progress at the end of the shift.

TCS
periodic chesks of TCS chillers; no refill needed.

On-call
from 5:05 UTC to 5:40 UTC remote intervention on SUSP; problem fixed
from 11:35 UTC to 12:20 UTC remote intervention on DAQ; problem fixed
from 12:24 UTC on site intervention of ISYS; in progress ...

Down-time
from 5:05 UTC to 5:40 UTC for SUSP on call intervention; ITF in Troubleshooting.
from 11:35 UTC to 12:20 UTC for DAQ on call intervention; ITF in Troubleshooting.
Images attached to this report
Comments to this report:
INJ team - 17:54 Tuesday 11 June 2019 (46081) Print this report

we found out that the increase of the temperature was not real, but just a probe failure. In fact, as you can see from the attached plot, the probe which is located on the pipe didn't report any variation, as well as the amplifier diodes temperatures.

We prepared another probe, which will be installed tomorrow morning.

Images attached to this comment
derossi - 17:32 Wednesday 12 June 2019 (46092) Print this report

Temperature probe back again (see attached plot).

Images attached to this comment
Search
×

Warning

×